OpenStudy (anonymous):

given the midpoint of CD is (2,-1) and point c is (-2,-4) what are the coordinates of D

OpenStudy (anonymous):

assume the coordinates of D are (x,y). for x: -2+x/2=2 x-2=4 x=6 for y: -4+y/2=2 y-4=4 y=8 Therefore,(x,y)=(6,8)

OpenStudy (anonymous):

i don't think thats right

OpenStudy (anonymous):

well,I think its correct.

OpenStudy (anonymous):

it multiple choice and thats not a choice

OpenStudy (anonymous):

what are the choices??

OpenStudy (anonymous):

(0,-2.5) (-2.5,0) (2,6) (6,2)

OpenStudy (anonymous):

i think its 6,2

OpenStudy (anonymous):

ohhh!!the midpoint is (2,-1)..I accidentally wrote -4+y/2=2,instead of writting -4+y/2=-1.Now calculate..

OpenStudy (anonymous):

:) thank you soo much you rock

OpenStudy (anonymous):

well,sorry for the mistake.:)

OpenStudy (anonymous):

its okay